# Nonlinear subspace clustering by functional link neural networks

### v1

@@ Abstract
Nonlinear subspace clustering based on a feed-forward neural network has been demonstrated to provide better clustering accuracy than some advanced subspace clustering algorithms. While this approach demonstrates impressive outcomes, it involves a balance between effectiveness and computational cost. In this study, we employ a functional link neural network to transform data samples into a nonlinear domain. Subsequently, we acquire a self-representation matrix through a learning mechanism that builds upon the mapped samples. As the functional link neural network is a single-layer neural network, our proposed method achieves high computational efficiency while ensuring desirable clustering performance. By incorporating the local similarity regularization to enhance the grouping effect, our proposed method further improves the quality of the clustering results. Additionally, we introduce a convex combination subspace clustering scheme, which combining a linear subspace clustering method with the functional link neural network subspace clustering approach. This combination approach allows for a dynamic balance between linear and nonlinear representations. Extensive experiments confirm the advancement of our methods. The source code will be released on https://lshi91.github.io/ soon.
@@ 1 Introduction
Subspace clustering involves the objective of identifying clusters within data points distributed across distinct subspaces in a high-dimensional space vidal2011subspace ; tang2024incomplete . Its practical applications encompass various scenarios, including image segmentation elhamifar2013sparse ; cao2023robust , motion segmentation liu2012robust ; favaro2011closed , face grouping liu2010robust ; qin2022maximum , character identification lu2018subspace ; patel2013latent , among others. The categorization of subspace clustering methods generally comprises four classes: algebraic approaches vidal2005generalized , iterative techniques ho2003clustering , statistical methodologies tipping1999mixtures ; rao2009motion , and methods based on spectral clustering li2017structured ; liu2014enhancing . Notably, spectral clustering-based techniques have garnered considerable attention due to their superior clustering performance. In essence, spectral clustering-based methods entail fundamental stages: 1) initial acquisition of a representation matrix to establish a similarity matrix derived from original data points, 2) subsequent execution of spectral decomposition and $k$-means clustering to derive the ultimate cluster assignments.
@@
Over the last decade, several subspace clustering techniques have emerged in literature, all aiming to uncover the underlying data subspace structure. The most popular spectral clustering algorithms are Sparse Subspace Clustering (SSC) elhamifar2013sparse and Low Rank Representation (LRR) liu2012robust . These methods enable sparse or minimum rank representation of data samples, respectively. Specifically, SSC operates under the assumption that each subspace can be expressed as a sparse linear combination of other points, while LRR aims to create a low-rank matrix that captures the overall data structure. To preserve both the global and local data structures, some researchers have explored incorporating constraints related to low rank and sparsity wang2016lrsr ; brbic2018ell_0 ; zhu2018low . The Frobenius norm is applied to the representation matrix in the Least Squares Regression (LSR) algorithm lu2012robust , with the goal of establishing the block diagonal characteristic within both inter-cluster and intra-cluster affinities. lu2013correlation merges SSC and LSR through the utilization of the trace Lasso technique. This adaptive approach enables the dynamic selection and clustering of correlated data. Lu et al. put forth an innovative regularization strategy relying on the block diagonal matrix, directly imposing the block diagonal structure onto the Block Diagonal Representation (BDR) lu2018subspace . Hu et al presented the grouping effect as the extent to which a representation matrix reflects the cluster structure of data points and developed the Smooth Representation Clustering (SMR) model hu2014smooth , which incorporates a smoothness regularizer to impose the grouping effect and a least squares error term to fit the data. Moreover, the author also introduced a novel affinity measure that leverages the grouping effect, which shows better performance than the conventional one. To deal with complex noise that often occurs in real-world scenarios, some robust subspace clustering algorithms based on maximum correntropy criterion shi2018improved ; shi2023efficient have been reported he2015robust ; guo2022correntropy ; zhang2019robust .
@@
Many existing techniques for subspace clustering rely on leveraging the linear connections between samples to learn the affinity matrix. In deed, in many practical scenarios, data samples often exhibit nonlinear relationships. The previous mentioned subspace clustering methods may experience performance degradation when dealing with such nonlinear data samples, thereby motivating the development of nonlinear subspace clustering methods. As a typical category of non-linear techniques, kernel-based subspace clustering methods have gained popularity. These methods possess the capability to perform a nonlinear transformation patel2014kernel ; yin2016kernel ; yang2019joint ; zhen2020kernel ; wang2021consensus . One such example is the Kernel Sparse Subspace Clustering (KSSC) patel2014kernel , which incorporates the kernel method into SSC, enabling the algorithm to find a nonlinear decision boundary that can separate the data more effectively. Kang et al introduced a comprehensive framework that unifies graph construction and kernel learning, harnessing the resemblance of the kernel matrix to enhance clustering performance kang2019low . Zhen et al enhanced the algorithm robustness and guaranteed the block diagonal property by truncating the trivial coefficients of kernel matrix zhen2020kernel . Multi-kernel learning methods have been used to find the optimal kernel combination for better clustering results zhang2021robust ; sun2021projective . In addition to the kernel method, Zhu et al employed a feed-forward neural network for executing nonlinear mapping and subsequently introduced the Nonlinear Subspace Clustering (NSC) technique zhu2018nonlinear . Deep Subspace Clustering (DSC) techniques that combine the advantages of deep neural networks and subspace clustering methods have drawn much attention peng2020deep , which introduce a self-expression layer between the encoder and the decoder to learn a similarity graph. However, subspace clustering methods based on neural network have limitations in computational cost when the layer number is large, and may also be sensitive to the setting of hyperparameters ji2017deep .
@@
To address the aforementioned challenge, this paper introduces an innovative subspace clustering technique that utilizes the functional link neural network (FLNN). This approach is termed Functional Link Neural Network Subspace Clustering (FLNNSC). FLNN is a single-layer artificial neural network that applies nonlinear transformations to the input features to strengthen the learning capability of the network patra2002nonlinear ; lu2019time ; yin2020combination . Compared with the conventional neural network, FLNN is able to achieve comparable function approximation performance with faster convergence and lower computational complexity. In practical scenarios, data often exhibit both linear and nonlinear structures. To fully exploit both the linear and nonlinear features of the data, we propose an extension of FLNNSC, introducing a convex combination subspace clustering (CCSC) method that simultaneously integrates linear and nonlinear representations. Our main contributions are summarized as follows:
@@
    We investigate a novel method known as FLNNSC for nonlinear subspace clustering. This method is highly effective in exploiting the nonlinear relations of data points. It also effectively captures local similarity through the grouping effect and mitigates overfitting by imposing a weight regularization mechanism. To solve the optimization problem for FLNNSC, we develop an iterative optimization algorithm.
@@
    By integrating linear and nonlinear representations, we propose a CCSC method, allowing for a comprehensive exploration and exploitation of the data’s diverse features.
@@
    By conducting experiments on widely recognized datasets, we confirm the superiority of our advanced subspace clustering approaches.
The organization of the paper is as follows. Section 2 offers a review of relevant subspace clustering methods. In Section 3, we introduce our novel subspace clustering approach. The evaluation of our methods’ performance across diverse datasets is conducted in Section 4. Conclusions are drawn in Section 5.
Notations. The main notations used in the paper are summarized in Table 1.
@@
2 Related work
In this section, we briefly review SSC elhamifar2013sparse and LRR liu2012robust . To facilitate our review, we first present some relevant definitions and notations. Let $\mathbf{X}=[\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{n}]\in\mathbb{R}^{d\times n}$ and $\mathbf{Z}=[\mathbf{z}_{1},\mathbf{z}_{2},\cdots,\mathbf{z}_{n}]\in\mathbb{R}^{n\times n}$ denote the data matrix and self-representation matrix, respectively, where $\mathbf{x}_{i}$ is the $i$th sample of $\mathbf{X}$, $d$ denotes the data dimension, and $n$ is the data number.
@@ 2 Related work
where the utilization of the $l_{1}$-norm serves the purpose of pursuing a sparse representation, the vector $diag(\cdot)\in\mathbb{R}^{n}$ is constructed from the diagonal entries of $\mathbf{Z}$, and the constraint $diag(\mathbf{Z})=\mathbf{0}$ is imposed to prevent the emergence of trivial solutions. Solving the optimization problem as stated in (1) can be achieved by implementing the Alternating Direction Method of Multipliers (ADMM) boyd2011distributed .
@@ 3 FLNN subspace clustering
This section first presents the basic principle of functional link neural network. Subsequently, we present the formulation of our novel model along with the procedures for optimization. The framework of our proposed FLNNSC method is shown in Figure 1.
@@
3.1 Functional link neural network
It is seen from the second panel of Figure 1 that the FLNN is a single-layer neural network, comprised of an input layer that is functionally expanded and an accompanying output layer. Notably, there is an absence of a hidden layer within this configuration. The input layer can be expanded by some nonlinear functions such as Chebyshev polynomials, Legendre polynomials, trigonometric functions, etc, which is useful to strengthen the learning capability of the network. For example, consider the $i$th element $x_{i}$ in a vector $\mathbf{x}$, an enhanced representation by using a second-order trigonometric polynomial is given by $\varphi(x_{i})=[x_{i},\sin(\pi x_{i}),\cos(\pi x_{i}),\sin(2\pi x_{i}),\cos(2\pi x_{i})]$. The FLNN has a simpler structure than the traditional feed-forward neural network, as it does not have any hidden layers. This reduces the computational complexity and the number of parameters that need to be trained.
@@ 3.2 Model formulation
where $\alpha$ is a balance parameter, the Laplacian matrix $\mathbf{L}$ is determined through the calculation $\mathbf{L}=\mathbf{D}-\mathbf{S}$ with $\mathbf{S}$ measuring the data similarity, and $\mathbf{D}$ being a diagonal matrix with its diagonal entries calculated by $D_{ii}=\sum^{n}_{j=1}S_{ij}$. A prevalent approach to construct $\mathbf{S}$ involves the use of a $k$ nearest neighbor ($k$-nn) graph, coupled with either a heat kernel or 00-$1$ weights hu2014smooth . The third component $J_{3}$ is defined as
@@ 3.3 Optimization
We execute iterative updates of $\mathbf{W}$ and $\mathbf{Z}$ until the objective function converges to a stable state. Subsequent to this, we attain the self-representation $\mathbf{Z}$ and proceed with the construction of the graph as outlined in hu2014smooth . Finally, we apply spectral clustering on the resulted graph $\mathbf{G}$. A comprehensive description of the FLNNSC approach is presented in Table 2
@@ 3.4 Computational complexity analysis
To initiate the convergence loop, we first compute the Laplacian matrix $\mathbf{L}$ in $O(n^{2})$ time using the formula $\mathbf{L}=\mathbf{D}-\mathbf{S}$. Within the loop, we compute $\varphi(\mathbf{x}_{i})$, update $\mathbf{W}$, and compute $\mathbf{H}$ for all data points at each iteration. Given that a second-order trigonometric polynomial expansion for $\varphi(\mathbf{x}_{i})$ is used, the dimensionality after the expansion is $\hat{d}=5d$, and the complexity of computing $\varphi(\mathbf{x}_{i})$ is $O(5d)$. The update of $\mathbf{W}$ involves two terms: ($\mathbf{h}_{i}-\mathbf{H}\mathbf{z}_{i})\odot\rho^{\prime}(\mathbf{W}\varphi(\mathbf{x}_{i}))\varphi^{T}(\mathbf{x}_{i})$ and $\beta\mathbf{W}$, with complexities of $O(125d^{3}+5dn^{2}+25d^{2}+5d)$ and $O(25d^{2})$, respectively. The complexity of computing $\mathbf{H}$ is $O(125d^{3}n)$. For $n$ data points, the overall complexity within the loop is $O(125d^{3}n^{2}+125d^{3}n+5dn^{3}+50d^{2}n+10dn)$. The update of $\mathbf{Z}$ is based on solving (15) by invoking the “lyap” function. This requires computing $\mathbf{H}^{T}\mathbf{H}$, which has a complexity of $O(25d^{2}n^{4})$, and calling the “lyap” function, which has a complexity of $O(2n^{3})$. Thus, the overall complexity of updating $\mathbf{Z}$ is $O(25d^{2}n^{4}+2n^{3})$. Assuming $t$ iterations, the final complexity corresponding to the convergence loop is $O(t(25d^{2}n^{4}+125d^{3}n^{2}+125d^{3}n+5dn^{3}+50d^{2}n+2n^{3}+10dn))$.

### v2

@@ Abstract
Nonlinear subspace clustering based on a feed-forward neural network has been demonstrated to provide better clustering accuracy than some advanced subspace clustering algorithms. While this approach demonstrates impressive outcomes, it involves a balance between effectiveness and computational cost. In this study, we employ a functional link neural network to transform data samples into a nonlinear domain. Subsequently, we acquire a self-representation matrix through a learning mechanism that builds upon the mapped samples. As the functional link neural network is a single-layer neural network, our proposed method achieves high computational efficiency while ensuring desirable clustering performance. By incorporating the local similarity regularization to enhance the grouping effect, our proposed method further improves the quality of the clustering results. We name our method as Functional Link Neural Network Subspace Clustering (FLNNSC). Furthermore, we propose a convex combination subspace clustering scheme that combines a linear subspace clustering method with the functional link neural network subspace clustering approach. This combination method is named as Convex Combination Subspace Clustering (CCSC), which allows for a dynamic balance between linear and nonlinear representations. Extensive experiments conducted on four widely used datasets, including Extended Yale B, USPS, COIL20, and ORL, demonstrate that both FLNNSC and CCSC outperform several state-of-art subspace clustering methods in terms of clustering accuracy. Our affinity graph experiments reveal that FLNNSC exhibits clear block diagonal structures. We provide recommendations for hyperparameters in FLNNSC by performing a parameter sensitivity analysis, and empirically verify the convergence of FLNNSC. Additionally, we show that FLNNSC has a lower computational cost compared to two high-performing methods.
@@ 1 Introduction
Subspace clustering involves the objective of identifying clusters within data points distributed across distinct subspaces in a high-dimensional space [1, 2]. Its practical applications encompass various scenarios, including image segmentation [3, 4], motion segmentation [5, 6], face grouping [7, 8], character identification [9, 10], among others. The categorization of subspace clustering methods generally comprises four classes: algebraic approaches [11], iterative techniques [12], statistical methodologies [13, 14], and methods based on spectral clustering [15, 16]. Notably, spectral clustering-based techniques have garnered considerable attention due to their superior clustering performance. In essence, spectral clustering-based methods entail fundamental stages: 1) initial acquisition of a representation matrix to establish a similarity matrix derived from original data points, 2) subsequent execution of spectral decomposition and $k$-means clustering to derive the ultimate cluster assignments.
@@
Over the last decade, several subspace clustering techniques have emerged in literature, all aiming to uncover the underlying data subspace structure. The most popular spectral clustering algorithms are Sparse Subspace Clustering (SSC) [3] and Low Rank Representation (LRR) [5]. These methods enable sparse or minimum rank representation of data samples, respectively. Specifically, SSC operates under the assumption that each subspace can be expressed as a sparse linear combination of other points, while LRR aims to create a low-rank matrix that captures the overall data structure. To preserve both the global and local data structures, some researchers have explored the integration of both the low rank and sparse regularizations. Typical works in this domain can be found in references [17] and [18]. Brbic et al. conducted an in-depth investigation of the $l_{0}$-norm on sparse regularization to address the over-penalization problem induced by the $l_{1}$-norm [19]. The Frobenius norm is applied to the representation matrix in the Least Squares Regression (LSR) algorithm [20], with the goal of establishing the block diagonal characteristic within both inter-cluster and intra-cluster affinities. [21] merges SSC and LSR through the utilization of the trace Lasso technique. This adaptive approach enables the dynamic selection and clustering of correlated data. Lu et al. put forth an innovative regularization strategy relying on the block diagonal matrix, directly imposing the block diagonal structure onto the Block Diagonal Representation (BDR) [9]. Hu et al presented the grouping effect as the extent to which a representation matrix reflects the cluster structure of data points and developed the Smooth Representation Clustering (SMR) model [22], which incorporates a smoothness regularizer to impose the grouping effect and a least squares error term to fit the data. Moreover, the author also introduced a novel affinity measure that leverages the grouping effect, which shows better performance than the conventional one. To deal with complex noise that often occurs in real-world scenarios, many subspace clustering methods takes into account robust cost functions to enhance the algorithm robustness. Correntropy, particularly the Maximum Correntropy Criterion (MCC), stands out as a popular robust criterion. It has found successful applications in signal processing and machine learning [23], owing to its ability to leverage high-order statistical information [24]. He et al. applied MCC directly to subspace clustering [25], while Guo et al. incorporated correntropy with low-rank regularization to enable robust motion segmentation [26]. Furthermore, Zhang et al. considered the schatten p-norm and correntropy together for kernel subspace clustering [27]. In addition to the aforementioned research, more recent interest on subspace clustering has shifted towards the multi-view field, involving aspects such as tensor technique [28], information bottleneck principle [29], latent representation [30].
@@
Many existing techniques for subspace clustering rely on leveraging the linear connections between samples to learn the affinity matrix. In deed, in many practical scenarios, data samples often exhibit nonlinear relationships. The previous mentioned subspace clustering methods may experience performance degradation when dealing with such nonlinear data samples, thereby motivating the development of nonlinear subspace clustering methods. As a typical category of non-linear techniques, kernel-based subspace clustering methods have gained popularity. One such example is the Kernel Sparse Subspace Clustering (KSSC) [31], which incorporates the kernel method into SSC, enabling the algorithm to find a nonlinear decision boundary that can separate the data more effectively. Yin et al. explored the self-expressive principle for symmetric positive definite matrices in the kernel subspace [32]. Yang et al considered the benefits of correntropy and block diagonal regularization when designing the kernel subspace clustering method [33]. Zhen et al enhanced the algorithm robustness and guaranteed the block diagonal property by truncating the trivial coefficients of kernel matrix [34]. Wang et al. integrated coefficient discriminant information and kernel subspace clustering into a unified process [35]. Kang et al introduced a comprehensive framework that unifies graph construction and kernel learning, harnessing the resemblance of the kernel matrix to enhance clustering performance [36]. Multi-kernel learning methods have been used to find the optimal kernel combination for better clustering results [37, 38]. In addition to the kernel method, Zhu et al employed a feed-forward neural network for executing nonlinear mapping and subsequently introduced the Nonlinear Subspace Clustering (NSC) technique [39]. Deep Subspace Clustering (DSC) techniques that combine the advantages of deep neural networks and subspace clustering methods have drawn much attention [40], which introduce a self-expression layer between the encoder and the decoder to learn a similarity graph. However, subspace clustering methods based on neural network have limitations in computational cost when the layer number is large, and may also be sensitive to the setting of hyperparameters [41]. For a clear comparison, we summarize the advantages and shortcomings of several typical methods in Table 1.
@@
To address the aforementioned challenge, this paper introduces an innovative subspace clustering technique that utilizes the functional link neural network (FLNN). This approach is termed Functional Link Neural Network Subspace Clustering (FLNNSC). FLNN is a single-layer artificial neural network that applies nonlinear transformations to the input features to strengthen the learning capability of the network [42]. This network has found applications in diverse areas such as filter design [43], resolution of the van der Pol-Duffing oscillator [44]. Compared with the conventional neural network, FLNN is able to achieve comparable function approximation performance with faster convergence and lower computational complexity. In practical scenarios, such as in the case of biological signals [45, 46], data often exhibit both linear and nonlinear structures. To fully exploit both the linear and nonlinear features of the data, we propose an extension of FLNNSC, introducing a convex combination subspace clustering (CCSC) method that simultaneously integrates linear and nonlinear representations. Our main contributions are summarized as follows:
@@
    Compared to existing subspace clustering methods, FLNNSC is computationally effective in exploiting the nonlinear relations among data points, thanks to the lightweight structural features of FLNN.
@@
    FLNNSC is capable of capturing local similarity through the grouping effect, as well as avoiding the overfitting issue by imposing a weight regularization mechanism.
@@
    Our subsequent proposed CCSC method allows for a more comprehensive exploration of the data’s diverse features, due to the integration of both linear and nonlinear representations.
The organization of the paper is as follows. Section 2 offers a review of relevant subspace clustering methods. In Section 3, we introduce our novel subspace clustering approach. The evaluation of our methods’ performance across diverse datasets is conducted in Section 4. Conclusions are drawn in Section 5. The main notations used in the paper are summarized in Table 2.
Table 1: Summary of the advantages and shortcomings of the reviewed methods.
  Methods Advantages Shortcomings
SSC [3] sparse data selection grouping effect/nonlinear data proficiency
LRR [5] grouping effect deficiency in subset selection/nonlinear data proficiency
LSR [20] grouping effect deficiency in subset selection/nonlinear data proficiency
CASS [21] balance between SSC and LSR nonlinear data proficiency
BDR [9] block diagonal property nonlinear data proficiency
SMR [22] grouping effect nonlinear data proficiency
KSSC [28] nonlinear mapping capability performance sensitive with kernel function
NSC [36] nonlinear mapping capability performance sensitive with hyperparameters
DSC [37] strong capabilities for diverse data high computational cost
@@
2 Preliminaries
In this section, we briefly review SSC [3] and LRR [5]. Then, we presents the basic principle of functional link neural network.
2.1 Review of SSC and LRR
To facilitate our review, we first present some relevant definitions and notations. Let $\mathbf{X}=[\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{n}]\in\mathbb{R}^{d\times n}$ and $\mathbf{Z}=[\mathbf{z}_{1},\mathbf{z}_{2},\cdots,\mathbf{z}_{n}]\in\mathbb{R}^{n\times n}$ denote the data matrix and self-representation matrix, respectively, where $\mathbf{x}_{i}$ is the $i$th sample of $\mathbf{X}$, $d$ denotes the data dimension, and $n$ is the data number.
@@ 2.1 Review of SSC and LRR
where the utilization of the $l_{1}$-norm serves the purpose of pursuing a sparse representation, the vector $diag(\cdot)\in\mathbb{R}^{n}$ is constructed from the diagonal entries of $\mathbf{Z}$, and the constraint $diag(\mathbf{Z})=\mathbf{0}$ is imposed to prevent the emergence of trivial solutions. Solving the optimization problem as stated in (1) can be achieved by implementing the Alternating Direction Method of Multipliers (ADMM) [47].
@@
2.2 Fundamentals of FLNN
It is seen from the second panel of Figure 1 that the FLNN is a single-layer neural network, comprised of an input layer that is functionally expanded and an accompanying output layer. Notably, there is an absence of a hidden layer within this configuration. The input layer can be expanded by some nonlinear functions [42], such as Chebyshev polynomials, Legendre polynomials, trigonometric functions, etc, which is useful to strengthen the learning capability of the network. For example, consider the $i$th element $x_{i}$ in a vector $\mathbf{x}$, an enhanced representation by using a second-order trigonometric polynomial is given by $\varphi(x_{i})=[x_{i},\sin(\pi x_{i}),\cos(\pi x_{i}),\sin(2\pi x_{i}),\cos(2\pi x_{i})]$. The FLNN has a simpler structure than the traditional feed-forward neural network, as it does not have any hidden layers. This reduces the computational complexity and the number of parameters that need to be trained.
@@ 3 FLNN subspace clustering
This section first introduces the formulation of our novel model. Then, we present the procedures for optimization. Finally, the computationally complexity of FLNNSC is analyzed. The framework of our proposed FLNNSC method is shown in Figure 1.
@@ 3.1 Model formulation
where $\alpha$ is a balance parameter, the Laplacian matrix $\mathbf{L}$ is determined through the calculation $\mathbf{L}=\mathbf{D}-\mathbf{S}$ with $\mathbf{S}$ measuring the data similarity, and $\mathbf{D}$ being a diagonal matrix with its diagonal entries calculated by $D_{ii}=\sum^{n}_{j=1}S_{ij}$. A prevalent approach to construct $\mathbf{S}$ is to use the $k$ nearest neighbor ($k$-nn) graph with a heat kernel or 0-1 weights [22]. Generally, 0-1 weighted $k$-nn graph is most commonly used due to its simplicity and desirable performance. Typically, a default value of $k$ is set to 4 [22], and we follow this setting in our experiments. The third component $J_{3}$ is defined as
@@ 3.2 Optimization
We execute iterative updates of $\mathbf{W}$ and $\mathbf{Z}$ until the objective function converges to a stable state. Subsequent to this, we attain the self-representation $\mathbf{Z}$ and proceed with the construction of the graph as outlined in [22]. Finally, we apply spectral clustering on the resulted graph $\mathbf{G}$. A comprehensive description of the FLNNSC approach is presented in Table 3
@@ 3.3 Computational complexity analysis
To initiate the convergence loop, we first compute the Laplacian matrix $\mathbf{L}$ in $O(n^{2})$ time using the formula $\mathbf{L}=\mathbf{D}-\mathbf{S}$. Within the loop, we compute $\varphi(\mathbf{x}_{i})$, update $\mathbf{W}$, and compute $\mathbf{H}$ for all data points at each iteration. Given that a second-order trigonometric polynomial expansion for $\varphi(\mathbf{x}_{i})$ is used, the dimensionality after the expansion is $\hat{d}=5d$, and the complexity of computing $\varphi(\mathbf{x}_{i})$ is $O(5d)$. The update of $\mathbf{W}$ involves two terms: ($\mathbf{h}_{i}-\mathbf{H}\mathbf{z}_{i})\odot\rho^{\prime}(\mathbf{W}\varphi(\mathbf{x}_{i}))\varphi^{T}(\mathbf{x}_{i})$ and $\beta\mathbf{W}$, with complexities of $O(125d^{3}+5dn^{2}+25d^{2}+5d)$ and $O(25d^{2})$, respectively. The complexity of computing $\mathbf{H}$ is $O(125d^{3}n)$. For $n$ data points, the overall complexity within the loop is $O(125d^{3}n^{2}+125d^{3}n+5dn^{3}+50d^{2}n+10dn)$. The update of $\mathbf{Z}$ is based on solving (15) by invoking the “lyap” function. This requires a complexity of $O(2n^{3})$. Assuming $t$ iterations, the final complexity corresponding to the convergence loop is $O(t(5dn^{3}+2n^{3}+125d^{3}n^{2}+125d^{3}n+50d^{2}n+10dn))$. When the condition $d<<n$ is satisfied, the model complexity can be approximated as $O(n^{3})$.
Upon reviewing the relevant literature, it becomes clear that several existing methods, including SSC, LRR, LRSC, and LSR, primarily differ in the norm constraint imposed on the regularization term. These methods share a similar structure of objective function, which leads to an approximate computational complexity of $O(n^{3})$ when $d<<n$. However, in comparing actual running time, SSC and LRR generally take longer than LSR. This is because problems induced by sparse and nuclear norms typically require more computational resources. From the corresponding literature of SMR, we know that its complexity is also $O(n^{3})$. The computational cost of NSC is heavily dependent on the choice of feed-forward neural networks, with more layers resulting in higher costs. The KSSC method actually has a slightly higher complexity than SSC due to the additional kernel mapping. The BDR method, which introduces a block diagonal regularization, has better learning efficiency than SSC or LRR in some data scenarios. Despite this, both KSSC and BDR can be considered as having a complexity of $O(n^{3})$ from a qualitative perspective.
Therefore, we can qualitatively argue that many subspace clustering methods have a complexity of $O(n^{3})$, including our proposed method. However, it is worth noting that they exhibit discrepancy in their actual execution time.